OpenStudy (osanseviero):

Which is the angle of this line? P1(-1,0) P2(0,-sqrt3) So the slope is -sqrt 3...

OpenStudy (osanseviero):

(It asks the angle related to the positive part of x

OpenStudy (osanseviero):

Tan -1 of sqrt of 3 is -60...should I add that to 180? or to 90? or what am I supposed to do?

OpenStudy (jdoe0001):

hmm are you doing vectors?

OpenStudy (osanseviero):

Yep

OpenStudy (jdoe0001):

\(\bf \textit{angle between two vectors }\\ \quad \\ cos(\theta)=\cfrac{u \cdot v}{||u||\ ||v||} \implies \cfrac{\text{dot product}}{\text{product of magnitudes}}\\ \quad \\ \theta = cos^{-1}\left(\cfrac{u \cdot v}{||u||\ ||v||}\right)\)

OpenStudy (osanseviero):

Oh...but this is just a line and it asks the angle with the x axis

OpenStudy (osanseviero):

It gives me two points of the lines...vectors are not necesary for this, I think

OpenStudy (jdoe0001):

well... haemm I'd think is 2 vectors in standard position, that is, their starting point is the origin

OpenStudy (jdoe0001):

though if you draw them, the angle is very visible, since notice the 0's one is a moving over one axis, the other over the other axis

OpenStudy (osanseviero):

But what do I need to do, then? :( I got this formula : Angle = tan-1 of sthe slope So that is -60, but this is part two of the graph, so I don't know what should I do now

OpenStudy (osanseviero):

Oh...I would add 90 to the 60...and 150 could do it

OpenStudy (jdoe0001):